OpenStudy (anonymous):

A circle in the fi rst quadrant with center on the curve y = 2x^2-27 is tangent to the y-axis and the line 4x = 3y. The radius of the circle is m/n where m and n are relatively prime positive integers. Find m + n.

OpenStudy (anonymous):

 like this right?

OpenStudy (anonymous):

this is not my answer but that is the first quadrant

OpenStudy (anonymous):

right?

OpenStudy (anonymous):

are you sure? The question says that the circle is only in the first quadrant, not all 4 of them...

OpenStudy (anonymous):

@HyperChemist , nop, the cercle is only in the first quadrant

OpenStudy (anonymous):

anyone any idea? :S

OpenStudy (mertsj):

I'm thinking. What class is this for?

OpenStudy (anonymous):

Well the center is where the equations y=2x^2-27 and 4x=3y. So basically (4/3)x=2x^2-27, and \[x=-7\sqrt{10}+2\] \[7\sqrt{10}+2\] Im a little stumped as to how to find the radius though

OpenStudy (anonymous):

Also the center would not be \[-7\sqrt{10}+2\] because the domain of the functions are confined to the first quadrant

OpenStudy (paxpolaris):